OpenStudy (anonymous):

Ray GL bisects angle FGM. Suppose angle FGL measures (6x +4)º while angle LGM measures (8x - 4)º. What is the measure of angle FGM?

OpenStudy (anonymous):

since ray GL bisects the angle, the two parts will b equal therefore 6x + 4 = 8x - 4 => 4 + 4 = 8x - 6x => 8 = 2x => x = 8/2 => x = 4 so angle FGL = 6x + 4 = 6(4) + 4 = 24 + 4 = 28 degree angle LGM = angle FGL = 28 degree so angle FGM = angle FGL + angle LGM = 28 degree + 28 degree = 56 degree

OpenStudy (anonymous):

Thanks!!!